OpenStudy (anonymous):

Isn't arcsec=cos? Since sec=1/cos the inverse of sec would be 1/(1/cos) which would be cos??

OpenStudy (anonymous):

@Clyde101 arcsecant is not the same as cosine. I think you are confusing the "-1" exponent of arcsecant with reciprocals when the "-1" in fact indicates the inverse function and not the reciprocal. To clear things up, observe this:\[\bf \sec(x) =\frac{1}{\cos(x)} \ \ and \ \ [\sec(x)]^{-1} =\frac{1}{\sec(x)}=\cos(x)\]\[\bf arcsec(x)=\sec^{-1}(x) \ne [\sec(x)]^{-1}=\cos(x)\]

OpenStudy (anonymous):

The -1 exponent outside the brackets indicates cos(x) and it also indicates that it's the reciprocal of sec(x). However the -1 exponent inside indicates the inverse function of sec(x) which is not the same as the reciprocal cos(x).

OpenStudy (anonymous):

OHHH ok, so then I need help with a problem...how would I solve y=arcsec(-1)

OpenStudy (anonymous):

So since arcsec is the inverse function of secant, we can re-write the question like this:\[\bf y=\sec^{-1}(-1) \implies \sec(y)=\frac{ 1 }{ \cos(y) }=-1\]Can you solve for it now?

OpenStudy (anonymous):

pi?

OpenStudy (anonymous):

For what value of y will 1/cos(y) result in -1. Notice that by re-arranging we get:\[\bf \cos(y)=-1\]Which 'y' would give us -1 as the answer?

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

And that's it. You got the answer.

OpenStudy (anonymous):

but cos is pi for this value as well...so why doesn't arcsec =cos always?

OpenStudy (anonymous):

No nonono they are not the same thing.

OpenStudy (anonymous):

on the unit circle...pi=cos-1

OpenStudy (anonymous):

Here, look at the graphs of both functions here: https://www.desmos.com/calculator/rkz053safx

OpenStudy (anonymous):

Do they look the same?

OpenStudy (anonymous):

not at all no...so why is the value the same on the unit circle in this instance?

OpenStudy (anonymous):

They are not. They have just 1 point of intersection at x = 1.108 and that's the only place they are ever the same.

OpenStudy (anonymous):

oh ok i get that

OpenStudy (anonymous):

Good. One thing to remember though, is that with inverse functions, the domain/range is usually limited in the cases of arccos(x), arcsin(x) and arcsec(x) and a few others. So when you do something like:\[\bf y=\cos^{-1}(x) \implies \cos(y)=x\]And you are given a value for x and you must get the value for y, you shouldn't forget that when you solve for y using cos(y) = x, that the value of y u find must be in the range of the inverse function. Cos(x) has a range of 1 to -1 but arccos(x) is the inverse, so the domain and range switch and to make sure that it remains a function (passes vertical line test), the domain gets restricted to -1 to 1 and range becomes 0 to pi.

OpenStudy (anonymous):

So when you solve for y for a certain inverse trig function, you must keep it's domain/range in mind. @Clyde101

OpenStudy (anonymous):

ok I will thanks that really helps alot! online pre-calc honors isn't easy on your own :)

OpenStudy (anonymous):

haha yep i will